OpenStudy (anonymous):

determine all points where this function is differentiable and where applicable give the formula for the derivative ... f(x+iy)=(x-y^2)+i(2xy-y)

OpenStudy (anonymous):

Should we differentiate this?!is it applicable?!

OpenStudy (anonymous):

what is \(i\) here ? \(\sqrt{-1}\) ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Sorry, but I think this question is beyond the capabilities of most of the people on here. Here's something that I hope can help you at least a bit. http://www.wolframalpha.com/input/?i=derivative+f%28x%2Biy%29%3D%28x-y^2%29%2Bi%282xy-y%29 Not sure why it has put it in the form it has but you should be able to rearrange the equation to get f'(x+iy).

OpenStudy (anonymous):

btw ... thanks ... :)

OpenStudy (anonymous):

are u familiar with Cauchy-Riemann equations for functions like \[f(z)=u(x,y)+i \ v(x,y)\]

OpenStudy (anonymous):

yes ... i know it

OpenStudy (anonymous):

well u have \(u(x,y)=x-y^2\) and \(v(x,y)=2xy-y\) Cauchy-Riemann equations states that for Derivability of \(f(z)\) at point \(z_0=x_0+i y_0\) u must have \(u_x=v_y\) and \(u_y=-v_x\)

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

so u just set up the equations for our particular case and see what happens and what are the possible values for \(x\) and \(y\) for Derivability

OpenStudy (anonymous):

yes ... so we will get -2y=-2y and 2x-1=1 ~> x=1 ...

OpenStudy (anonymous):

now should we differentiate it?!or it is not applicable?!

OpenStudy (anonymous):

well Cauchy-Riemann equations are necessary for Derivability but not sufficient lol i cant explain my english is bad @Neemo

OpenStudy (anonymous):

:D ... thanks .

OpenStudy (anonymous):

Cauchy-Riemann equations are necessary and sufficient conditions for complex differentiation once we assume that its real and imaginary parts are differentiable real functions of two variables.

OpenStudy (anonymous):

A function f(z) = u(x, y) + i v(x, y), where z = x + i y, has a complex derivative f′(z) if and only if its real and imaginary parts are continuously differentiable and satisfy the Cauchy–Riemann equations : (mentionned by @mukushla)

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

@parto note that \(u\) and \(v\) are continuously differentiable and according to what u got satisfy the Cauchy–Riemann equations for \(x=1\)

OpenStudy (anonymous):

 Mapping of complex functions am I right?